Small samples and large population variances imply that the estimate for the mean will be relatively poor. Whether or not it will result in an underestimate or overestimate depends on the distribution: with a symmetric distribution the two outcomes are equally likely.
1/2 * 1/3 = 1/6. You can get 1/6 cup by using 2 tablespoons plus 2 teaspoons.
a firearm is a gun that fires bullets they're normally all have firing pins which hit the back of the shell causing a spark in the shell which ignites the gun powder and sends the bullet shooting of at over the speed of sounds.
When two full packs of cards are used what are the chances of drawing an ace?
Since every full pack has four aces, there would be eight aces in two packs. You would have a very good chance of drawing one or more aces.
15
19
only 1 because the wqhen you rollthe 2nd 6 you have to move the pawn u took out
You have a 3.125% chance of not rolling an even number, because each time you roll a die you have a 50% chance of not rolling an even number, but each additional time you roll a die your chances of not rolling an even number the formula changes from 3/6 to 3/12 because the possibilities double but your chances of not rolling an even remain the same so eventually we end up with 3/96 because of rolling the die 6 times.
How can study more than 15 hours per day.?
Well it will be hard or is should say very hard but it is not impossible.
like straight after coming for school then sit straight for study and create a atmosphere that you are in a exam and study for 3 hours not even getting up for drinking water then take a break and think what you would achieve by studying this much during break and again study.
Hope it will be useful.
Who is on the head and tail side of the penny?
The head of a penny is obviously the queens head, and the tail is that thing on the other side of the penny.
There are 72 permutations of two dice and one coin.
What do you mean by step and touch potential and it's standard value?
Step Potential is the potential Difference between two points on the earth's surface, separated by distance of one pace, that will be assumed to be one meter in the direction of maximum potential gradient (IS 3043) Touch Potential is the potential difference between a grounded metallic structure and a point on the earth's surface separated by a distance equal to the normal maximum horizontal reach, approximately one meter. (IS 3043)
What is the probable explanation for referred pain?
Nerves carry the signal to other body parts so your brain interprets the pain from being from that distal part also.
Decision making theory is used to determine the values and other issues, including uncertainties, that relate to the decision being made. It is then determined if the decision is a rational and wise decision to be made.
Determine the equation what line slope is -2 and y intercept equals -3?
y=mx+b; m=slope & b=y intercept.
Therefore y=-2x+(-3) or y=-2x-3.
What is probably being misunderstood by a person who says but that's only a scientific theory?
Nearly 100%. Someone who claims something is "just a theory" has already proven they lack an understanding of scientific definitions. This usually means they lack an understanding of the subject which they refered to as just a theory.
How many different ways can all of the letters of a nine letter word be arranged?
Factorial 9 (9!), ie 9 x 8 x 7 x 6 x 5 x 4 x 3 x 2 x 1 different ways, that's 362880!
What is the probability of rolling a die 120 times and getting 5 or 6?
It is 1 - pr(no 5 nor 6 in 120 rolls)
= 1 - (2/3)120
= 1 - 1.113*1057
which is so close to 1 as to make no difference.
It is 0.99...98887 which starts off with a string of 56 nines.
Assuming:
Number of possible rolls = 36 (6x6)
28 rolls satisfy "at least one die shows a 4 or the sum of the dice is greater than 5":
(1, 4), (1, 5), (1, 6),
(2, 4), (2, 5), (2, 6),
(3, 3), (3, 4), (3, 5), (3, 6),
(4, 1), (4, 2), (4, 3), (4, 4), (4, 5), (4, 6),
(5, 1), (5, 2), (5, 3), (5, 4), (5, 5), (5, 6),
(6, 1), (6, 2), (6, 3), (6, 4), (6, 5), (6, 6),
The probability that at least on die shows a 4 or that the sum of the dice is greater than 5 is 28/36, or 7/9, or approx. 77.8%
In a six sided die what is the possibility of rolling 2 odd numbers in a row?
There are 3 odd (1,3,5) and 3 even (2,4,6) sides on a 6-sided die.
When rolling a 6-sided die, each face has an equal chance of coming up (assuming a perfect die).
Because of this, the chance of rolling an odd number once is 50%.
The second roll you again have a 50% chance to roll an odd number, which means that the chance to roll two in a roll is (50% * 50%) 25%.
1/2*1/2 = 1/4 =
25%
In other words, the chance of rolling an odd number is 1/2 to the Nth power, where N is the number of rolls.
Rolling a dice 10 times means a chance of (1/2)n10 or
1/2 *1/2 *1/2 *1/2 *1/2 *1/2 *1/2 *1/2 *1/2 *1/2 =
1/1024 =
0,097%